# A Prompt-Guided Spatio-Temporal Transformer Model for National-Wide Nuclear Radiation Forecasting

### v1

@@ Abstract.
Nuclear radiation (NR), which refers to the energy emitted from atomic nuclei during decay, poses substantial risks to human health and environmental safety. Accurate forecasting of nuclear radiation levels is crucial for informed decision-making by both individuals and governments. However, this task is challenging due to the imbalanced distribution of monitoring stations over a wide spatial range and the non-stationary radiation variation patterns. In this study, we introduce NRFormer, an innovative framework tailored for national-wide prediction of nuclear radiation variations. By integrating a non-stationary temporal attention module, an imbalance-aware spatial attention module, and a radiation propagation prompting module, NRFormer collectively captures complex spatio-temporal dynamics of nuclear radiation. Extensive experiments on two real-world datasets demonstrate the superiority of our proposed framework against seven baselines. This research not only enhances the accuracy and reliability in nuclear radiation forecasting but also contributes to advancing emergency response strategies and monitoring systems, thereby safeguarding environmental and public health.
Spatio-Temporal Modeling, Transformer, Radiation Forecasting
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The Web of Things (WoT)(Hui et al., 2022) has seen significant growth and implementation in recent years, enabling the collection of various types of spatio-temporal data, including check-in data, satellite and street-view imagery, and sensor readings. These bring about many new emerging web-based applications e.g., urban region profiling (Yan et al., 2024), spatio-temporal modeling (Zheng et al., 2023; Yang et al., 2024), and web mining (Kang et al., 2024; Ye et al., 2024; Liu et al., 2024a; Lai et al., 2024). A key web-based application is nuclear radiation forecasting (see Appendix A), which can help people understand future radiation levels in their surroundings, enabling precise risk assessments and early disaster warnings, thereby safeguarding public health and environmental safety.
In this paper, we study the national-wide nuclear radiation forecasting problem, which involves collaborative forecasting across distributed monitoring stations throughout the country. Figure 1 presents an illustrative example of radiation forecasting in Japan. The dispersion of nuclear radiation is primarily influenced by the global circulation system (i.e., atmospheric and ocean circulation) (Qiao et al., 2011), manifesting intricate spatio-temporal patterns. Therefore, an effective forecasting model must be capable of comprehending these spatio-temporal correlations from geo-distributed stations. Recent advances in Spatio-Temporal Graph Neural Networks (STGNNs) (Jin et al., 2023; Wu et al., 2020; Shao et al., 2022c; Lan et al., 2022; Zhang et al., 2022; Shao et al., 2022a), which combine Graph Neural Networks (GNNs) with sequential models (e.g., Recurrent Neural Networks), have shown superiority in various applications such as traffic forecasting (Wu et al., 2019) and air quality prediction (Han et al., 2023). Nonetheless, existing STGNNs often fall short in capturing the complex dynamics of nuclear radiation, primarily due to the following three major challenges.
Figure 2. Empirical study of the imbalanced spatial distribution problem. (1) Full station distribution of our constructed Japan-1D dataset. (2) Evenly sampled station distribution (from Japan-1D dataset). (3) The forecasting performance of GWN on Japan datasets, where GWN-w/o SPM is the model variant that removes the spatial propagation module.
@@
Challenge 2: Imbalanced spatial distribution. Radiation monitoring stations are often deployed unevenly, with dense clusters near nuclear facilities and sparse distribution in other regions. Typical STGNNs capture spatial correlations by representing these stations as graph nodes and propagating information between adjacent nodes. However, such an approach is less effective in scenarios with spatial imbalances, as nodes in densely populated regions tend to be easily overwhelmed by abundant information, while sparsely connected nodes in other areas fail to perceive sufficient spatial context for prediction. For instance, we illustrate our insights using GWN (Wu et al., 2019), a widely adopted spatio-temporal model. As reported in Figure 2, we observe that removing the spatial propagation module of GWN (i.e., GWN-w/o SPM) diminished forecasting effectiveness on a balanced subset, whereas the opposite results are observed when using the full dataset. Such results indicate that the utility of the graph learning module is reduced in datasets with significant spatial imbalances. Consequently, how to capture diversified dependencies from uneven spatial distribution is also a critical challenge.
@@
  • We investigate the national-wide nuclear radiation forecasting problem. To the best of our knowledge, this is the first work to collectively predict radiation on a national scale.
  • We propose NRFormer, a model tailor-designed to tackle unique challenges in nuclear radiation forecasting. We have also built a web-based platform that provides interactive insights into radiation forecasting.
  • We construct two large-scale nuclear radiation datasets named Japan-4H and Japan-1D, encompassing three years of nuclear radiation data from 3,841 monitoring stations across Japan. We release the dataset to facilitate future research in this area. Extensive experiments validate the effectiveness of NRFormer against 11 baselines.
@@ 2. Preliminaries
In this section, we introduce some important definitions and formally define the nuclear radiation forecasting problem.
@@
3. Dataset Construction and Analysis
In this section, we introduce the datasets used for NRFormer with a preliminary data analysis.
3.1. Data Preprocessing
The datasets include nuclear radiation data and meteorological data collected from the entire country of Japan. Table 1 summarizes the statistics of the radiation and meteorological datasets.
@@ 3.1. Data Preprocessing
Nuclear radiation data. The original nuclear radiation data are collected from Japan’s Nuclear Regulation Authority (NRA), consisting of radiation measurements at 10-minute intervals from more than 4,000 monitoring stations, spanning from March 17th, 2021, to January 1st, 2024. We exclude radiation monitoring stations with over 30 days of missing data, obtaining a more reliable dataset encompassing 3,841 stations. To study the nuclear radiation forecasting from both long-term and short-term perspectives, we preprocess the original dataset to create two experimental datasets: (1) Japan-4H consists of 4-hour averaged radiation values with 6,121 time steps, and (2) Japan-1D includes daily averaged radiation values with 1,021 time steps.
@@
Meteorological data. Considering the significant influence of meteorological conditions on the spread of radiation, we collect Integrated Surface Dataset (ISD) from the National Oceanic and Atmospheric Administration (NOAA). The ISD dataset includes hourly surface meteorological observations from over 35,000 weather stations worldwide, covering a variety of parameters like pressure, temperature, dew point, winds, clouds, precipitation, ocean waves, tides, and so on. By using the geographic coordinates, we identify 235 weather stations in Japan and match them to corresponding radiation monitoring stations. We extract key meteorological variables, including wind speed, wind direction, air temperature, and dew point, as contextual features for radiation forecasting task.
@@
3.2. Data Analysis
To help understand the radiation distribution, we conduct a primary analysis of the radiation and meteorological data, as shown in Figure 3. First, we visualize the spatial distribution of radiation stations and weather stations in Figure 3(a) and  3(c), respectively. We observe that radiation stations exhibit a highly skewed spatial distribution, i.e., densely deployed in high-risk areas close to nuclear facilities. Next, we plot the temporal distribution of average radiation levels across all stations in Figure 3(b). As can be seen, since March 17, 2021, radiation levels have shown a clear downward trend, with January being the lowest of the year, which may be due to specific meteorological conditions during this period.
We further provide a detailed spatial and temporal distribution for adjacent prefectures. As illustrated in Figure 4, it is obvious that the radiation level varies over space and time non-linearly and significantly. The radiation patterns of spatially adjacent regions may not always be more similar than distant regions due to complex external factors, e.g., unpredictable human activities. Such complexities pose a significant challenge for the effective modeling of spatio-temporal correlations, especially for those monitoring stations in dense clusters.
In addition, we conduct an empirical analysis to explore the relationships between meteorology and the radiation level. Specifically, Figure 3(d) shows the temporal distribution of average daily values for four meteorological variables: wind speed, wind direction, air temperature, and dew point. We observe that the patterns of temperature and dew point closely resemble the daily radiation distribution shown in Figure 3(b). Although we do not observe similar results in wind speed and direction, they may implicitly correlate with nuclear radiation by influencing its propagation. Thus, we incorporate these variables to enhance forecasting.
[…]
Figure 5. The framework overview of NRFormer.
[…]
4. The Proposed Method
Overview. The framework of NRFormer is shown in Figure 5, which is comprised of three major modules: (1) The Non-stationary Temporal Attention: it integrates instance-wise normalization into point-wise temporal attention to capture the non-stationary dynamics of radiation time series. (2) The Imbalance-aware Spatial Attention: it models both global and local spatial dependencies to selectively propagate spatial knowledge for nodes with extremely dense or sparse neighborhoods. (3) The Radiation Propagation Prompting: it utilizes contextual features to provide a contextual prompt for the imbalanced-aware spatial attention module, enhancing prediction with essential environmental and spatial insights. Finally, the output layer fuses the embeddings from temporal and spatial attention, and then denormalizes them to produce the final prediction results. Next, we present each component in detail.
4.1. Non-stationary Temporal Attention
4.1.1. Non-stationary normalization
where $\gamma_{i}$ and $\beta_{i}$ are learnable parameters corresponding to the $i$-th station, and $\epsilon$ is a small value that ensures numerical stability. Intuitively, the normalized sequences possess more stable statistical properties, which largely reduce the difficulty of capturing temporal dynamics within the non-stationary data. Subsequently, we process $\hat{\mathbf{x}}_{i}$ through an embedding layer:
(4) $\mathbf{z}_{t}^{i}=Embed(\hat{\mathbf{x}}_{i}),$
where $\mathbf{z}_{t}^{i}\in\mathbb{R}^{P\times D}$ represents the point-wise time series embedding of station $i$ that will be fed into subsequent temporal attention module, $D$ is the hidden dimension of each time step.
4.1.2. Point-wise temporal attention
where $\mathbf{W}_{t}^{Q},\mathbf{W}_{t}^{K},\mathbf{W}_{t}^{V}\in\mathbb{R}^{D\times D}$ are learnable parameters shared across different stations and time steps. Afterward, we compute the attention score $A_{t}$ to measure the relationships among any two time steps
(6) $\mathbf{A}_{t}=\frac{\mathbf{Q}_{t}{\mathbf{K}_{t}}^{\top}}{\sqrt{D}},$
where $A_{t}\in\mathbb{R}^{P\times P}$ and $\sqrt{D}$ is the scaling factor.
We normalize each row of $\mathbf{A}_{t}$ with a Softmax function, and obtain the updated representation vector as follows
(7) $\mathbf{z}_{t}^{i,(l)}=Softmax(\mathbf{A}_{t})\mathbf{V}_{t}.$
To enhance the effectiveness of temporal dependency modeling, we further equip multi-head attention and residual connection technique for each layer, followed by layer normalization (Vaswani et al., 2017). We build the temporal attention module by stacking $L_{t}$ such layers, and denote the output of the $L_{t}$-th layer as $\mathbf{z}_{t}^{i,(L_{t})}$.
Finally, we flatten $\mathbf{z}_{t}^{i,(L_{t})}$ into a 1D vector and map the flattened version into a hidden representation
(8) $\mathbf{z}_{t}^{i}=Flatten(\mathbf{z}_{t}^{i,(L_{t})})\mathbf{W}^{r},$
where $\mathbf{W}^{r}$ is learnable parameters. Let $\mathbf{Z}_{t}\in\mathbb{R}^{N\times D}$ denotes the hidden representations of all the nodes (i.e., stations), where the $i$-th row vector of $\mathbf{Z}_{t}$ is $\mathbf{z}_{t}^{i}$.
4.2. Imbalance-aware Spatial Attention
As previously mentioned, the nodes in dense clusters have the risk of information overload (i.e., over-smoothing) and sparsely connected nodes have the risk of information insufficient (i.e., under-smoothing) in spatial dependency modeling, leading to sub-optimal performance. To alleviate this problem, we develop the imbalance-aware spatial attention, which includes two parts: (1) a global spatial attention module to enrich the spatial neighborhood of orphan nodes, and (2) a local spatial attention module to reduce the influence of noisy neighborhood for densely connected nodes.
Following previous works (Vaswani et al., 2017; Lai et al., 2023), we add positional encoding into $\mathbf{Z}_{s}$ to preserve critical spatial information of each node:
(9) $\mathbf{Z}_{s}^{(0)}=\mathbf{Z}_{t}+\mathbf{P}_{e},$
where $\mathbf{P}_{e}\in\mathbb{R}^{N\times D}$ is a learnable positional embedding matrix. The resulting matrix $\mathbf{Z}_{s}^{(0)}$ will be fed into the following spatial attention block for spatial dependency modeling.
4.2.1. Global spatial attention
In this part, we employ the global spatial attention mechanism to identify latent neighborhoods for sparsely-connected nodes. Specifically, we first compute the attention score between arbitrary nodes:
(11) $\mathbf{A}_{g}=\frac{\mathbf{Q}_{g}\cdot\mathbf{K}_{g}^{\top}}{\sqrt{D}},$
where $\mathbf{A}_{g}\in\mathbb{R}^{N\times N}$ is the attention score matrix that encodes the all-pair node relationships. Then, we can propagate node signals based on the normalized $\mathbf{A}_{g}$, defined as
(12) $\mathbf{H}_{g}=Softmax(\mathbf{A}_{g})\cdot\mathbf{V}_{g}.$
Next, we further adopt a feed-forward layer to enhance the model expressive capabilities
(13) $\mathbf{H}^{\prime}_{g}=(ReLU(\mathbf{H}_{g}\cdot\mathbf{W}_{1}+\mathbf{b}_{1}))\cdot\mathbf{W}_{2}+\mathbf{b}_{2}.$
where $\mathbf{W}_{1},\mathbf{W}_{2}$ are the learnable matrices and $\mathbf{b}_{1},\mathbf{b}_{2}$ are the bias parameters. In summary, the global spatial attention can provide additional spatial information for nodes with scarce neighborhoods, which reduces the risk of under-smoothing phenomena.
4.2.2. Local spatial attention
In practice, the global spatial attention may introduce massive noises for densely-connected nodes via all-pair message passing, which exacerbates the over-smoothing issue. Therefore, we devise the local spatial attention by adding a graph constraint to filter those neighbors that are far away from the target nodes. Likewise, we first compute the attention score matrix $\mathbf{A}_{l}$ and the key matrix $\mathbf{V}_{l}$ via Equations 10 and 11. Suppose $\mathbf{A}$ denotes the truncated adjacency matrix built from geographical distance, we define a masked matrix $\mathbf{A}_{mask}$ based on $\mathbf{A}$ and $\mathbf{A}_{l}$:
where $\mathbf{A}_{mask}[i,j]$ is set to $\infty$- ∞ if the distance between node $i$ and $i$ exceeds a threshold. Afterwards, we perform spatial message passing under local constraint:
(15) $\mathbf{H}_{l}=Softmax(\mathbf{A}_{mask})\cdot\mathbf{V}_{l}.$
The output $\mathbf{H}_{l}$ is then fed into a feed-forward layer,
where $\mathbf{W}_{3},\mathbf{W}_{4}$ are the learnable matrices and $\mathbf{b}_{3},\mathbf{b}_{4}$ are the bias parameters. The local spatial message passing operator can restrict the spatial receptive field of densely connected nodes and enable the adaptive selection of useful neighboring nodes, which can avoid potential over-smoothing risk. We alternately stack $L_{s}$ global and local spatial attention layers for spatial dependency modeling. The final output is denoted as $\mathbf{Z}_{s}^{(L_{s})}$.
@@ 4.2.2. Local spatial attention
Data Models 6th 9th 12th 24th sudden change
@@
4.3. Radiation Propagation Prompting
Meteorological prompt. Meteorological conditions significantly influence the spatial propagation of nuclear radiation. For example, strong wind can carry radioactive material over long distances, affecting areas far from the contamination source. Therefore, for each radiation monitoring station, we retrieve the observations of its nearest meteorological station as prompt features, denoted as $\mathbf{M}\in\mathbb{R}^{N\times T\times C_{m}}$, where $C_{m}$ is the number of meteorological features. We flatten the last two dimensions of $\mathbf{M}$ into a 2D matrix $\mathbf{M}^{\prime}\in\mathbb{R}^{N\times C_{o}}$, where $C_{o}=T\times C_{m}$. Similarly, we transform the raw features into a latent feature representation through an MLP
(17) $\mathbf{E}_{m}=MLP_{m}(\mathbf{M}^{\prime}).$
After that, we fuse the above two prompt feature vectors $\mathbf{E}_{p}=\mathbf{E}_{l}||\mathbf{E}_{m}$, where $\mathbf{E}_{p}$ is the contextual prompt embedding that will be injected into the imbalanced-aware spatial attention as addition clues. Concretely, we can rewrite Equation 10 as

### v2

# NRFormer: Nationwide Nuclear Radiation Forecasting with Spatio-Temporal Transformer

@@ Abstract.
Nuclear radiation, which refers to the energy emitted from atomic nuclei during decay, poses significant risks to human health and environmental safety. Recently, advancements in monitoring technology have facilitated the effective recording of nuclear radiation levels and related factors, such as weather conditions. The abundance of monitoring data enables the development of accurate and reliable nuclear radiation forecasting models, which play a crucial role in informing decision-making for individuals and governments. However, this task is challenging due to the imbalanced distribution of monitoring stations over a wide spatial range and the non-stationary radiation variation patterns. In this study, we introduce NRFormer, a novel framework tailored for the nationwide prediction of nuclear radiation variations. By integrating a non-stationary temporal attention module, an imbalance-aware spatial attention module, and a radiation propagation prompting module, NRFormer collectively captures complex spatio-temporal dynamics of nuclear radiation. Extensive experiments on two real-world datasets demonstrate the superiority of our proposed framework against 11 baselines. NRFormer has been deployed online to provide 1–24-day nuclear radiation forecasts, empowering individuals and governments with timely, data-driven decisions for emergency response and public safety. Our framework is designed for general applicability and can be readily adapted for deployment in other regions. The deployed system is available at https://NRFormer.github.io and the dataset and code of the predictive model are available at https://github.com/usail-hkust/NRFormer.
nuclear radiation forecasting system, spatio-temporal modeling, imbalance-aware transformer
@@ 1. Introduction
In this paper, we study the national-wide nuclear radiation forecasting problem, which involves collaborative forecasting across distributed monitoring stations throughout the country. Figure 1 shows an illustrative example of radiation forecasting in Japan. The dispersion of nuclear radiation is primarily influenced by the global circulation system (i.e., atmospheric and ocean circulation) (Qiao et al., 2011), manifesting intricate spatio-temporal patterns. Therefore, an effective forecasting model must be capable of comprehending these spatio-temporal correlations from geo-distributed stations. Recent advances in Spatio-Temporal Graph Neural Networks (STGNNs) (Jin et al., 2023; Wu et al., 2020; Shao et al., 2022c; Lan et al., 2022; Zhang et al., 2022; Shao et al., 2022a), which combine Graph Neural Networks (GNNs) with sequential models (e.g., RNN), have shown superiority in various applications such as traffic forecasting (Wu et al., 2019; Bai et al., 2020) and air quality prediction (Han et al., 2023). Nonetheless, existing STGNNs often fall short in capturing the complex dynamics of nuclear radiation, primarily due to the following three major challenges.
Figure 2. Empirical study of the imbalanced spatial distribution problem. (1) Full station distribution of our constructed Japan-1D dataset. (2) Evenly sampled station distribution (from Japan-1D dataset). (3) The forecasting performance of GWN on Japan datasets, where GWN-w/o SPM is the model variant that removes the spatial propagation module (SPM).
@@
Challenge 2: Imbalanced spatial distribution. Radiation monitoring stations are often deployed unevenly, with dense clusters near nuclear facilities and sparse distribution in other regions. Typical STGNNs capture spatial correlations by representing these stations as graph nodes and propagating information between adjacent nodes. However, such an approach is less effective in scenarios with spatial imbalances, as nodes in densely populated regions tend to be easily overwhelmed by abundant information, while sparsely connected nodes in other areas fail to perceive sufficient spatial context for prediction. For instance, we illustrate our insights using GWN (Wu et al., 2019), a widely adopted spatio-temporal model. As reported in Figure 2, we observe that removing the spatial propagation module of GWN (i.e., GWN-w/o SPM) diminished forecasting effectiveness on a balanced subset, whereas the opposite results are observed when using the full dataset. Such results indicate that the utility of the graph learning module is compromised, leading to reduced accuracy in datasets with significant spatial imbalances. Consequently, how to capture diversified dependencies from uneven spatial distribution is also a critical challenge.
@@
  • We investigate the nationwide nuclear radiation forecasting problem. To the best of our knowledge, this is the first work to collectively predict radiation on a national scale.
  • We propose NRFormer, a predictive model specifically designed to address distinct challenges in nuclear radiation forecasting. By integrating a non-stationary temporal attention module, an imbalance-aware spatial attention module, and a radiation propagation prompting module, NRFormer effectively captures complex spatio-temporal dynamics inherent to nuclear radiation phenomena.
  • We construct two large-scale nuclear radiation datasets named Japan-4H and Japan-1D, encompassing three years of nuclear radiation data from 3,841 monitoring stations across Japan. We release the dataset  [^1] https://github.com/usail-hkust/NRFormer to facilitate future research in this area. Extensive experiments validate the effectiveness of NRFormer against 11 baselines.
  • We deploy a web-based, real-time radiation forecasting system  [^2] https://NRFormer.github.io, providing 1–24-day predictions and rich radiation data analysis, while offering interactive insights into spatial-temporal radiation patterns.
@@ 2. Preliminaries
3. System Overview
3.1. System Architecture
Figure 4 illustrates the architecture of our deployed radiation forecasting system. The Data Sources consist of publicly accessible web services that provide real-time radiation and meteorological data. The Data Collector systematically acquires time-series data from external sources through web-service APIs or automated web-scraping techniques. During the Data Processing and Data Analysis stages, raw data are cleaned and stored in a cloud-based database. The Predictive Model leverages this refined dataset to generate forecasts of radiation levels for each radiation monitoring station over the next 1 to 24 days. These forecasts are stored in the cloud database and retrieved by the Web Service, which provides access to forecasting results and enables interactive visualization and analysis of radiation data through dynamic web interfaces.
Figure 4. Architecture of deployed forecasting system.
[…]
3.2. Data Collecting and Processing
@@
Nuclear radiation data. The nuclear radiation data are continuously collected from Japan’s Nuclear Regulation Authority  [^3] https://www.erms.nsr.go.jp/nra-ramis-webg/, consisting of 10-minute interval radiation measurements from over 4,000 monitoring stations. Stations with more than 30 days of missing data were excluded, resulting in a refined dataset of 3,841 stations. To analyze radiation forecasting from both long-term and short-term perspectives, the original data were preprocessed to generate two datasets: (1) Japan-4H, comprising 4-hour averaged values, and (2) Japan-1D, containing daily averaged values.
@@
Meteorological data. Considering the significant influence of meteorological conditions on the spread of radiation, we continuously collect Integrated Surface Dataset (ISD) from the National Oceanic and Atmospheric Administration (NOAA)  [^4] https://www.ncei.noaa.gov/metadata/geoportal/rest/metadata/item/gov.noaa.ncdc:C00532/html. The ISD dataset includes hourly surface meteorological observations from over 35,000 weather stations worldwide, covering a variety of parameters like pressure, temperature, dew point, winds, and so on. By using the geographic coordinates, we identify 235 weather stations in Japan and match them to corresponding radiation monitoring stations. We extract key meteorological variables, including wind speed, wind direction, air temperature, and dew point, as contextual features for radiation forecasting task.
[…]
3.3. Data Analysis
To investigate radiation distribution patterns, we perform a preliminary analysis of radiation and meteorological data, as illustrated in Figure 3. First, Figures 3(a) and 3(c) visualize the spatial distributions of radiation and weather stations, respectively. The analysis reveals a highly skewed spatial distribution of radiation stations, i.e., , dense clustering near high-risk nuclear facilities. Figure 3(b) displays the temporal distribution of average radiation levels, showing a pronounced downward trend beginning March 17, 2021, with January exhibiting the lowest annual levels—a pattern potentially linked to seasonal meteorological conditions.
We further analyze spatial and temporal distributions across adjacent prefectures in Figure 5, which demonstrates significant non-linear variations in radiation levels across both dimensions. Notably, radiation patterns in geographically proximate regions are not necessarily more similar than those in distant areas, likely due to external complexities such as unpredictable human activities. These dynamics complicate the modeling of spatio-temporal correlations, particularly for densely clustered monitoring stations.
Additionally, we empirically examine relationships between meteorological factors and radiation levels. Figure 3(d) illustrates the temporal distributions of four meteorological variables: wind speed, wind direction, air temperature, and dew point. Temperature and dew point exhibit daily patterns strongly correlated with radiation trends in Figure 3(b). While wind parameters lack explicit correlations, their potential implicit influence on radiation propagation justifies their inclusion. This analysis confirms statistically significant relationships between meteorological conditions and radiation levels, with their integration substantially improving accuracy.
3.4. User Interfaces
We developed a web-based real-time forecasting system to provide nationwide radiation level predictions across Japan. The deployed system is presented in Figure 10 and 11 in the Appendix and available at https://NRFormer.github.io.
This deployed system enables users to monitor temporal changes in radiation levels. As shown in Figure 10, each monitoring station is represented by a blue marker on the map. Selecting a station triggers a pop-up chart that depicts forecasted radiation levels for subsequent days, offering insights into projected radiation trends. These visualizations assist stakeholders in environmental and public health risk assessment and mitigation planning. Additionally, Figure 10 features a heatmap displaying daily radiation forecasts across Japan. The heatmap highlights geospatial variations in radiation levels, revealing patterns and dispersion dynamics over time. This enables policymakers to identify regional disparities and temporal progression of radiation exposure risks.
As illustrated in Figure 11, the system provides multifaceted radiation analysis through multiple visualization modalities. This system allows both individuals and government agencies to examine radiation data from diverse perspectives, fostering a holistic comprehension of nuclear risks and supporting data-driven decision-making processes.
Figure 6. The framework overview of NRFormer.
[…]
4. The Predictive Model
Overview. The framework of NRFormer is shown in Figure 6, which is comprised of three modules: (1) The Non-stationary Temporal Attention: it integrates instance-wise normalization into point-wise temporal attention to capture the non-stationary dynamics of radiation. (2) The Imbalance-aware Spatial Attention: it models both macro-scale and proximity-constrained spatial dependencies to selectively propagate spatial knowledge for nodes with extremely dense or sparse neighborhoods. (3) The Radiation Propagation Prompting: it utilizes contextual features to provide a contextual prompt for the imbalanced-aware spatial attention module, enhancing prediction with essential environmental and spatial insights. Finally, the output layer fuses the embeddings from temporal and spatial attention, and then denormalizes them to produce the final prediction results. Next, we present each component in detail.
4.1. Non-stationary Temporal Attention
4.1.1. Non-stationary normalization
where $\gamma_{i}$ and $\beta_{i}$ are learnable parameters corresponding to the $i$-th station, and $\epsilon$ is a small value that ensures numerical stability. Intuitively, the normalized sequences possess more stable statistical properties, which largely reduce the difficulty of capturing temporal dynamics within the non-stationary data. Subsequently, we process $\hat{\mathbf{x}}_{i}$ through an embedding layer: $\mathbf{z}_{t}^{i}=Embed(\hat{\mathbf{x}}_{i})$, where $\mathbf{z}_{t}^{i}\in\mathbb{R}^{P\times D}$ represents the point-wise time series embedding of station $i$ that will be fed into subsequent temporal attention module, $D$ is the hidden dimension of each time step.
4.1.2. Point-wise temporal attention
where $\mathbf{W}_{t}^{Q},\mathbf{W}_{t}^{K},\mathbf{W}_{t}^{V}\in\mathbb{R}^{D\times D}$ are learnable parameters shared across different stations and time steps. Afterward, we compute attention score $A_{t}$ to measure relationships among any two time steps, $\mathbf{A}_{t}=\frac{\mathbf{Q}_{t}{\mathbf{K}_{t}}^{\top}}{\sqrt{D}}$, where $A_{t}\in\mathbb{R}^{P\times P}$ and $\sqrt{D}$ is the scaling factor.
We normalize each row of $\mathbf{A}_{t}$ with a Softmax function, and obtain the updated representation vector as follows $\mathbf{z}_{t}^{i,(l)}=Softmax(\mathbf{A}_{t})\mathbf{V}_{t}$. To enhance the effectiveness of temporal dependency modeling, we further equip multi-head attention and residual connection technique for each layer, followed by layer normalization (Vaswani et al., 2017). We build the temporal attention module by stacking $L_{t}$ such layers, and denote the output of the $L_{t}$-th layer as $\mathbf{z}_{t}^{i,(L_{t})}$.
Finally, we flatten $\mathbf{z}_{t}^{i,(L_{t})}$ into a 1D vector and map the flattened version into a hidden representation, $\mathbf{z}_{t}^{i}=Flatten(\mathbf{z}_{t}^{i,(L_{t})})\mathbf{W}^{r}$, where $\mathbf{W}^{r}$ is learnable parameters. Let $\mathbf{Z}_{t}\in\mathbb{R}^{N\times D}$ denotes the hidden representations of all the nodes (i.e., stations), where the $i$-th row vector of $\mathbf{Z}_{t}$ is $\mathbf{z}_{t}^{i}$.
4.2. Imbalance-aware Spatial Attention
As previously mentioned, the nodes in dense clusters have the risk of information overload (i.e., over-smoothing) and sparsely connected nodes have the risk of information insufficient (i.e., under-smoothing) in spatial dependency modeling, leading to sub-optimal performance. To alleviate this problem, we develop the imbalance-aware spatial attention, which includes two parts: (1) a macro-scale spatial correlation modeling module to enrich the spatial neighborhood of orphan nodes, and (2) a proximity-constrained spatial modeling module to reduce the influence of noisy neighborhood for densely connected nodes.
Following previous works (Vaswani et al., 2017; Lai et al., 2023), we add positional encoding into $\mathbf{Z}_{s}$ to preserve critical spatial information of each node: $\mathbf{Z}_{s}^{(0)}=\mathbf{Z}_{t}+\mathbf{P}_{e}$, where $\mathbf{P}_{e}\in\mathbb{R}^{N\times D}$ is a learnable positional embedding matrix. The resulting matrix $\mathbf{Z}_{s}^{(0)}$ will be fed into the following spatial attention block for spatial dependency modeling.
4.2.1. Macro-Scale Spatial Correlation Modeling
In this part, we employ the macro-scale spatial attention mechanism to identify latent neighborhoods for sparsely-connected nodes. Specifically, we first compute the attention score between arbitrary nodes:
(6) $\mathbf{A}_{g}=\frac{\mathbf{Q}_{g}\cdot\mathbf{K}_{g}^{\top}}{\sqrt{D}},$
where $\mathbf{A}_{g}\in\mathbb{R}^{N\times N}$ is the attention score matrix that encodes the all-pair node relationships. Then, we can propagate node signals based on the normalized $\mathbf{A}_{g}$, defined as $\mathbf{H}_{g}=Softmax(\mathbf{A}_{g})\cdot\mathbf{V}_{g}$. Next, we further adopt a feed-forward layer to enhance the model expressive capabilities, $\mathbf{H}^{\prime}_{g}=(ReLU(\mathbf{H}_{g}\cdot\mathbf{W}_{1}+\mathbf{b}_{1}))\cdot\mathbf{W}_{2}+\mathbf{b}_{2}$, where $\mathbf{W}_{1},\mathbf{W}_{2}$ are the learnable matrices and $\mathbf{b}_{1},\mathbf{b}_{2}$ are the bias parameters. In summary, macro-scale spatial attention can provide additional spatial information for nodes with scarce neighborhoods, which reduces the risk of under-smoothing phenomena.
4.2.2. Proximity-Constrained Spatial Correlation Modeling
In practice, the macro-scale spatial attention may introduce massive noises for densely-connected nodes via all-pair message passing, which exacerbates the over-smoothing issue. Therefore, we devise proximity spatial attention by adding a graph constraint to filter those neighbors that are far away from the target nodes. Likewise, we first compute the attention score matrix $\mathbf{A}_{l}$ and the key matrix $\mathbf{V}_{l}$ via Equations 5 and 6. Suppose $\mathbf{A}$ denotes the truncated adjacency matrix built from geographical distance, we define a masked matrix $\mathbf{A}_{mask}$ based on $\mathbf{A}$ and $\mathbf{A}_{l}$:
where $\mathbf{A}_{mask}[i,j]$ is set to $\infty$- ∞ if the distance between node $i$ and $i$ exceeds a threshold. Afterwards, we perform spatial message passing under proximity constraint: $\mathbf{H}_{l}=Softmax(\mathbf{A}_{mask})\cdot\mathbf{V}_{l}$. The output $\mathbf{H}_{l}$ is then fed into a feed-forward layer,
where $\mathbf{W}_{3},\mathbf{W}_{4}$ are the learnable matrices and $\mathbf{b}_{3},\mathbf{b}_{4}$ are the bias parameters. The proximity spatial message passing operator can restrict the spatial receptive field of densely connected nodes and enable the adaptive selection of useful neighboring nodes, which can avoid potential over-smoothing risk. We alternately stack $L_{s}$ macro-scale and proximity spatial attention layers for spatial dependency modeling. The final output is denoted as $\mathbf{Z}_{s}^{(L_{s})}$.
4.3. Radiation Propagation Prompting
Meteorological prompt. Meteorological conditions significantly influence the spatial propagation of nuclear radiation. For example, strong wind can carry radioactive material over long distances, affecting areas far from the contamination source. Therefore, for each radiation monitoring station, we retrieve the observations of its nearest meteorological station as prompt features, denoted as $\mathbf{M}\in\mathbb{R}^{N\times T\times C_{m}}$, where $C_{m}$ is the number of meteorological features. We flatten the last two dimensions of $\mathbf{M}$ into a 2D matrix $\mathbf{M}^{\prime}\in\mathbb{R}^{N\times C_{o}}$, where $C_{o}=T\times C_{m}$. Similarly, we transform the raw features into a latent feature representation through an MLP: $\mathbf{E}_{m}=MLP_{m}(\mathbf{M}^{\prime})$. After that, we fuse the above two prompt feature vectors $\mathbf{E}_{p}=\mathbf{E}_{l}||\mathbf{E}_{m}$, where $\mathbf{E}_{p}$ is the contextual prompt embedding that will be injected into the imbalanced-aware spatial attention as addition clues. Concretely, we can rewrite Equation 5 as
@@ 4.3. Radiation Propagation Prompting
Data Models 6th (24 hours / 6 days) 9th (36 hours / 9 days) 12th (48 hours / 12 days) 24th (96 hours / 24 days) sudden change